OpenStudy (anonymous):

solve x+2y=-4 -3x+2y=4

OpenStudy (anonymous):

x+2y=-4 -3x+2y=4 x=-2 y=-1

OpenStudy (anonymous):

(x+2y)*(-3)=-4*(-3)

OpenStudy (anonymous):

-3x-6y=12

OpenStudy (anonymous):

-3x-6y=12 -3x+2y=4 - - - 0+-8y=8 y=-1 now put y=-1 -3x-6*(-1)=12 x=-2

OpenStudy (anonymous):

lets verify put x=-2, Y=-1 -3(-2)-6(-1)=12 12=12 LHs=RHS